OpenStudy (anonymous):

If x is an integer greater than 1 and y=x+1 divided by x, which of the following must be true? I. y does not equal x. II. y is an integer. III. xy is greater than x squared

OpenStudy (anonymous):

You know, I is true and II is false. III is up to you. Do you know why?

OpenStudy (anonymous):

well

OpenStudy (anonymous):

xy will always be greater than x sq. because y is always x+1/x

OpenStudy (anonymous):

Well, consider this. Let x be 10. y = 10 + 1 / 10 = 11/10 xy = 10· 11 / 10 = 11 and x² = 121. As you can see, III is not true.

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

I mean x² = 100. Sorry.

OpenStudy (anonymous):

the answer booklet says I and III are true

OpenStudy (anonymous):

you still there?

OpenStudy (anonymous):

That's impossible... Are you sure?

OpenStudy (anonymous):

what do you say, son goku?

hartnn (hartnn):

@micahwood50 y=10+1/10 = 101/10

OpenStudy (anonymous):

Or do you actually mean \[y = x + \frac{1}{x}\]

OpenStudy (anonymous):

x+1/x

OpenStudy (anonymous):

x and one is not in numerator together, right?

hartnn (hartnn):

so III is always true

hartnn (hartnn):

ay y>x xy>x^2

OpenStudy (anonymous):

How come 10+1/10 = 101/10?

hartnn (hartnn):

*as

OpenStudy (anonymous):

you combine digits together?

hartnn (hartnn):

10+1/10 = (10*10+1)/10 = 101/10

hartnn (hartnn):

y= x+1/x so y>x xy>x^2

OpenStudy (anonymous):

Well, 101 / 10 = 10.1 and x² = 100 so...

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

good job, guys

OpenStudy (anonymous):

Let me show you. \[xy \rightarrow x\left(\frac{x+1}{x}\right) = x + 1\] x + 1 > x² Or if y = x + (1/x), then: \[xy \rightarrow x\left( x + \frac{1}{x} \right) = x^2 + 1\] x² + 1 > x² In this case, you mean x + (1/x) so yeah, III is true. Sorry for confusion.

OpenStudy (anonymous):

good job, guys3

OpenStudy (anonymous):

thank you guys for all of the help

OpenStudy (anonymous):

these questions are in prep for the PSATs

OpenStudy (anonymous):

x + 1 < x² ***

OpenStudy (anonymous):

You're welcome. Glad we helped.